# Tackling Negative Transfer on Graphs

### v1

@@ 4.3 Ablation Study
Transfer without Fine-tuning. In previous settings, we fine-tune the model classifier to evaluate the quality of knowledge encoded in backbones. Following Liu et al. (2023a), we take a further step by directly employing the pre-trained model on target graph without fine-tuning. Table 6 presents the transfer learning performance on Citation and Arxiv benchmarks. Note that we adopt another domain adaptation setting (Degree) from Gui et al. (2022). It is obvious that our proposed SP layer significantly improves the transfer learning performance by enhancing the quality of the encoder.
@@
The reported results are from Liu et al. (2023a). $*{}^{*}$ indicates the results of our implementations based on the official code.
@@
Transfer with Fully Fine-tuning. We also analyze the transfer learning results when fine-tuning the whole model. Table 5 shows the results that even if GNN-SP only fine-tunes the last layer, it still outperforms standard GNN with fully fine-tuning. If we fine-tune the whole model of GNN-SP, the model performance can be further improved, especially on large-scaled Arxiv dataset.
Pooling Methods. Apart from adopting multiple backbones, we also evaluate model performance with different pooling functions, including MEAN, ATTN, MAX, and GCN. Figure 5 presents the experimental results over Citation network with GCN backbone. We observe the basic MEAN pooling outperforms complicated GCN and ATTN that adaptively determine the pooling weights $w_{ij}$. It is might because the complicated methods introduce extra inductive bias with unexpected noises.
@@ 5 Related Works
Pre-training with Target Graph. Researchers developed methods to explicitly align source and target graphs during pre-training, ensuring that $P_{S}(\mathcal{Z})=P_{T}(\mathcal{Z})$. For example, Zhang et al. (2019); Dai et al. (2022b) adopt adversarial learning Ganin et al. (2016) to train domain-invariant encoder, and following UDAGCN Wu et al. (2020) incorporates attention to further enhance expressiveness. Alternatively, one can employ regularizers Zhu et al. (2021a, 2023); Shi et al. (2023), such as MMD and CMD, to constrain the discrepancy between the source and target. To facilitate this process, new graph-specific discrepancy metrics are proposed, including tree mover distance Chuang and Jegelka (2022), subtree discrepancy Wu et al. (2023a), and spectral regularizer You et al. (2023). Additionally, Liu et al. (2023a) emphasizes the most relevant instances in the source to better match distributions of the source and target. While these methods effectively reduce the distribution shift between source and target, the target graph may be unavailable during pre-training in many real-world scenarios.
@@
Pre-training without Target Graph. Existing works aim to train GNNs that can be transferred to unseen target graphs. For example, EERM Wu et al. (2022b) and following works Chen et al. (2022); Wu et al. (2022c, 2023b); Yu et al. (2023) utilize causal learning to develop environment-invariant encoder. GTrans Jin et al. (2023) transforms the target graph at test-time to align the source and target. Moreover, various studies employ augmentation to enhance the robustness of encoder against permutations Verma et al. (2021); Wang et al. (2021); Liu et al. (2022); Han et al. (2022); Liu et al. (2023b); Guo et al. (2023) or apply disentangle learning to extract domain-invariant semantics Ma et al. (2019); Liu et al. (2020). To understand the transferability of GNNs, Ruiz et al. (2020); Levie et al. (2021); Bevilacqua et al. (2021); Cao et al. (2023) interpret graphs as the combination of graphons, while Han et al. (2021); Zhu et al. (2021b); Sun et al. (2023); Qiu et al. (2020) adopt self-supervised learning to identify transferable structures. Despite these efforts to enhance transferability, they cannot well address the negative transfer issue. To this end, we systematically analyze why the negative transfer happens and provide insights to solve this issue.
@@ 6 Conclusion
In this paper, we explore the negative transfer in GNNs and introduce Subgraph Pooling, a simple yet effective method, to mitigate the issue. Our method transfers the subgraph-level knowledge to reduce the discrepancy between the source and target graphs, and is applicable for any GNN backbone without introducing extra parameters. We provide a theoretical analysis to demonstrate how the model works and conduct extensive experiments to evaluate its superiority under various transfer learning settings.
@@ Appendix A Proof
Proof. To demonstrate this theorem, we analyze the distance between the subgraph embeddings $\|\mathbf{h}_{u}-\mathbf{h}_{v}\|$ in terms of the node embeddings:
@@
For this term, under Assumption 1, it is sufficiently small to be disregarded. Combining these analyses, we derive $\|\mathbf{h}_{u}-\mathbf{h}_{v}\|\leq\|\mathbf{z}_{u}-\mathbf{z}_{v}\|-\Delta$. While $\Delta$ may not always be positive, it is essential for understanding the impact of SP. To further substantiate this, we present two corollaries.
@@
Proof. To establish that $\Delta\geq 0$, we need to demonstrate that:
$n\|\mathbf{z}_{u}-\mathbf{z}_{v}\|$ $\geq\frac{m-n}{m+1}\|\mathbf{z}_{v}\|)$ (21)
@@
Proof. Assuming the target graph is richer than the source, indicated by $|\mathcal{N}_{s}(v)|\geq|\mathcal{N}_{s}(u)|$, it is necessary to validate that $\|\mathbf{z}_{u}-\mathbf{z}_{v}\|\geq\|\mathbf{z}_{v}\|$ to maintain the inequality. This validation becomes crucial in the extreme case where $n\to 0$ and $m\to\infty$, leading to $\lim_{n\to 0,m\to\infty}\frac{(m-n)}{n(m+1)}=1$. We must demonstrate that $\frac{\|\mathbf{z}_{u}-\mathbf{z}_{v}\|}{\|\mathbf{z}_{v}\|}\geq\frac{(m-n)}{n(m+1)}$ is valid under these circumstances, implying $\frac{\|\mathbf{z}_{u}-\mathbf{z}_{v}\|}{\|\mathbf{z}_{v}\|}\geq 1$. Therefore, when the source is sparse, it may lack sufficient information for effective transfer.
@@
where $\cos\theta=\mathbf{z}_{x}^{T}\mathbf{z}_{y}/\|\mathbf{z}_{x}\|\|\mathbf{z}_{y}\|$. According to Definition 2, this inequality is satisfied when $\lambda\geq 2$.
@@ Appendix C Hyper-parameters
Table 8: Hyper-parameters.
@@ D.2 More Analysis
Varying Training Ratio. The transfer learning performance highly relies on the number of training instances available on the target. To evaluate the sensitivity to varying training sample sizes, we consider four training ratios $\{0.1\%,0.5\%,1\%,10\%\}$, as shown in Table 2 in the main body. Our propose GNN-SP outperforms all baselines. We observe that transfer learning models achieve better performance than the model trained from scratch (No Transfer) with extremely limited labels. This is because the knowledge gained from the source provides beneficial regularization during the fine-tuning process. However, when the training ratio reaches $q=10\%$, standard GNNs acquire sufficient knowledge to surpass other baselines, with the exception of our GNN-SP.

### v2

@@ 4.3 Ablation Study
Transfer without Fine-tuning.
Following Liu et al. (2023a), we take a further step by directly employing the pre-trained model on target graph without fine-tuning. Table 6 presents the transfer learning performance on Citation and Arxiv. We also adopt another domain adaptation setting (Degree) from Gui et al. (2022). It is obvious that our proposed SP layer significantly improves the transfer learning performance by enhancing the quality of the encoder.
@@
The reported results are from Liu et al. (2023a). indicates the results of our implementations based on the official code.
@@
Transfer with Fully Fine-tuning.
We present transfer learning performance by fine-tuning the whole model in Table 5. We observe even if GNN-SP only fine-tunes the last layer, it still outperforms standard GNN with full fine-tuning. If our GNN-SP is fully fine-tuned, the model performance can be further improved, especially on large-scaled Arxiv.
Pooling Methods.
We evaluate model performance with different pooling functions, including MEAN, ATTN, MAX, and GCN. Figure 5 presents the experimental results over Citation network with GCN backbone. We observe the basic MEAN pooling outperforms complicated GCN and ATTN that adaptively determine the pooling weights $w_{ij}$. It is might because the complicated methods introduce extra inductive bias with unexpected noises.
@@ 5 Related Works
Pre-training with Target Graph.
Researchers developed methods to explicitly align source and target graphs during pre-training, ensuring $P_{S}(\mathcal{Z})=P_{T}(\mathcal{Z})$. For example, Zhang et al. (2019); Dai et al. (2022b) adopt adversarial learning Ganin et al. (2016) to train domain-invariant encoder, and following UDAGCN Wu et al. (2020) incorporates attention to further enhance expressiveness. Alternatively, one can employ regularizers Zhu et al. (2021a, 2023); Shi et al. (2023), such as MMD and CMD, to constrain the discrepancy between the source and target. To facilitate this process, new graph-specific discrepancy metrics are proposed, including tree mover distance Chuang and Jegelka (2022), subtree discrepancy Wu et al. (2023a), and spectral regularizer You et al. (2023). Additionally, Liu et al. (2023a) emphasizes the most relevant instances in the source to better match distributions of the source and target. While these methods effectively reduce the distribution shift between source and target, the target graph may be unavailable during pre-training in many real-world scenarios.
@@
Pre-training without Target Graph.
Existing works aim to train GNNs transferable to unseen graphs. For example, EERM Wu et al. (2022b) and following works Chen et al. (2022); Wu et al. (2022c, 2023b); Yu et al. (2023) utilize causal learning to develop environment-invariant encoder. GTrans Jin et al. (2023) transforms the target graph at test-time to align the source and target. Moreover, various studies employ augmentation to enhance the robustness of encoder against permutations Verma et al. (2021); Wang et al. (2021); Liu et al. (2022); Han et al. (2022); Liu et al. (2023b); Guo et al. (2023) or apply disentangle learning to extract domain-invariant semantics Ma et al. (2019); Liu et al. (2020). To understand the transferability of GNNs, Ruiz et al. (2020); Levie et al. (2021); Cao et al. (2023) interpret graphs as the combination of graphons, while Han et al. (2021); Zhu et al. (2021b); Sun et al. (2023); Qiu et al. (2020) adopt self-supervised learning to identify transferable structures. Despite these efforts to enhance transferability, they cannot well address the negative transfer issue. To this end, we systematically analyze why the negative transfer happens and provide insights to solve this issue.
@@ 6 Conclusion
In this paper, we analyze the negative transfer in GNNs and introduce Subgraph Pooling, a simple yet effective method, to mitigate the issue. Our method transfers the subgraph-level knowledge to reduce the discrepancy between the source and target graphs, and is applicable for any GNN backbone without introducing extra parameters. We provide a theoretical analysis to demonstrate how the model works and conduct extensive experiments to evaluate its superiority under various transfer learning settings.
Acknowledgements
This work was partially supported by the NSF under grants IIS-2321504, IIS-2334193, IIS-2340346, IIS-2203262, IIS-2217239, CNS-2203261, and CMMI-2146076. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Appendix A Proof
Proof.
To demonstrate this theorem, we analyze the distance between the subgraph embeddings $\|\mathbf{h}_{u}-\mathbf{h}_{v}\|$ in terms of the node embeddings:
@@
For this term, under Assumption 1, it is sufficiently small to be disregarded. Combining these analyses, we derive $\|\mathbf{h}_{u}-\mathbf{h}_{v}\|\leq\|\mathbf{z}_{u}-\mathbf{z}_{v}\|-\Delta$. While $\Delta$ may not always be positive, it is essential for understanding the impact of SP. To further substantiate this, we present two corollaries. ∎
@@
Proof.
To establish that $\Delta\geq 0$, we need to demonstrate that:
$n\|\mathbf{z}_{u}-\mathbf{z}_{v}\|$ $\geq\frac{m-n}{m+1}\|\mathbf{z}_{v}\|$ (21)
@@
Proof.
Assuming the target graph is richer than the source, indicated by $|\mathcal{N}_{s}(v)|\geq|\mathcal{N}_{s}(u)|$, it is necessary to validate that $\|\mathbf{z}_{u}-\mathbf{z}_{v}\|\geq\|\mathbf{z}_{v}\|$ to maintain the inequality. This validation becomes crucial in the extreme case where $n\to 0$ and $m\to\infty$, leading to $\lim_{n\to 0,m\to\infty}\frac{(m-n)}{n(m+1)}=1$. We must demonstrate that $\frac{\|\mathbf{z}_{u}-\mathbf{z}_{v}\|}{\|\mathbf{z}_{v}\|}\geq\frac{(m-n)}{n(m+1)}$ is valid under these circumstances, implying $\frac{\|\mathbf{z}_{u}-\mathbf{z}_{v}\|}{\|\mathbf{z}_{v}\|}\geq 1$. Therefore, when the source is sparse, it may lack sufficient information for effective transfer.
@@
where $\cos\theta=\mathbf{z}_{x}^{T}\mathbf{z}_{y}/\|\mathbf{z}_{x}\|\|\mathbf{z}_{y}\|$. According to Definition 2, this inequality is satisfied when $\lambda\geq 2$. ∎
@@ Appendix C Hyper-parameters
Table 8: Hyper-parameters for our proposed Subgraph Pooling and Subgraph Pooling++.
@@ D.2 More Analysis
Varying Training Ratio. The transfer learning performance highly relies on the number of training instances available on the target. To evaluate the sensitivity to varying training sample sizes, we consider four training ratios $\{0.1\%,0.5\%,1\%,10\%\}$, as shown in Table 2 in the main body. Our proposed GNN-SP outperforms all baselines. We observe that transfer learning models achieve better performance than the model trained from scratch (No Transfer) with extremely limited labels. This is because the knowledge gained from the source provides beneficial regularization during the fine-tuning process. However, when the training ratio reaches $q=10\%$, standard GNNs acquire sufficient knowledge to surpass other baselines, with the exception of our GNN-SP.